OpenStudy (calculusxy):

MEDAL!!! Weights of different materials were taken for three lengths: 6 inches, 12 inches, and 18 inches. Material 1: 15, 30, 60 Material 2: 13, 26, 39 Material 3: 31, 62, 93 Material 4: 14, 28, 42 Which of the materials' weight does NOT have a proportional relationship with length? A. Material 1 B. Material 2 C. Material 3 D. Material 4

OpenStudy (calculusxy):

@Michele_Laino

OpenStudy (calculusxy):

@Nnesha

OpenStudy (anonymous):

Material 1: 15, 30, 60 is a difference of +15 then +30 Material 2: 13, 26, 39 is a difference of +13, +13 Material 3: 31, 62, 93 is a difference of +31, +31 Material 4: 14, 28, 42 is a difference of +14, +14 For material 1: 6 / 15 = 12 / 30 = 18 / 60 0.4 = 0.4 = 0.3 0.4 does not equal 0.3. Therefore, Material 1 will not work.

OpenStudy (anonymous):

So that means that Material 1 is your answer

OpenStudy (calculusxy):

Okay thank you! I have another question.

OpenStudy (anonymous):

Your welcome and sure!